a man is 65 years old and his grandson is 20. how many years ago was the grandfather six times as old as the grandson?
Answer by orca(409)

Suppose that x years ago the grandfather was six times as old as the grandson.
x years ago, their ages were as follows:
Grandfather: 65 - x,
Grandson: 20 - x.
As then the grandfather was six times as old as the grandson, we can set up an equation:
65 - x = 6(20 - x)
Solving for x, we obtain:
65 - x = 120 - 6x
5x = 55
x = 11
So 11 years ago the grandfather was six times as old as the grandson.